# Interpretable factorization of clinical questionnaires to identify latent factors of psychopathology

### 12 Dec 2023 v1

@@ 3.2 Solving the optimization problem
with high probability. The full proof of Proposition 2 is provided in Appendix 6.4. The two propositions, combined, show that our factorization can capture the true latent structure of the data, under the right conditions. The first is a linear combination of factors being a good approximation, which is the case for questionnaires. The second is having a robust estimator of $k$, discussed next.
Choice of number of factors  For each $\beta$, we choose the number of factors $k$ using blockwise-cross-validation (BCV). Given a matrix $M$, for each $k$, we shuffle the rows and columns of $M$ and subdivide it into $b_{r}\times b_{c}$ blocks. These blocks are split into 10 folds and we repeatedly omit blocks in a fold, factorize the remainder, impute the omitted blocks via matrix completion and compute the error [^1] Appropriate weighting is multiplied to the error if number of blocks in the last fold is less than others. of that imputation. We choose $k$ with the lowest average error. This procedure can adapt to the distribution of confounds $C$ by stratified splitting. We compared this with other approaches for choosing $k$, for ICQF and other methods, over synthetic data, and report the results in Section 4.1.
@@ 4.1 Experiments on synthetic questionnaire data
We examined the effectiveness of BCV and other algorithms on estimating the number of latent factors in a synthetic dataset, for ICQF against $\ell_{1}$-regularized NMF ($\ell_{1}$-NMF) [15] and factor analysis with promax rotation (FA-promax) [32] as factors can be correlated. Both ICQF and $\ell_{1}$-NMF were initialized with NNDSVD [10], and the sparsity ($\beta=1\mathrm{e}{-1}$) and stopping criterion (relative iteration convergence tolerance $\epsilon<1\mathrm{e}{-3})$ for fairness. The estimation method for FA was minimum residual.
We generated a synthetic questionnaire with $k^{*}=10$ factors. To generate a $200\times 10$ latent factor matrix $W$ (Figure 1 left), we first define a matrix $D$, a binary matrix in which columns are correlated with a step-like pattern. Each column is present in isolation for 20 participants, and in tandem with another for 10 more, to synthesize correlation between factors. Then an entry of $W[i,j]$ is defined as
@@
Each factor had an associated loading vector – answer pattern – over 100 questions ($[0,100]$ range). The resulting $100\times 10$ loading matrix $Q$ , shown in Figure 1 (center), is defined to be
@@
We then create a noiseless data matrix $M_{clean}:=\min(0,\max(WQ^{T},100))$, and add noise by
@@
where $e$ follows a discrete probability distribution with $P(e=1)=\delta,P(e=0)=1-\delta$. This yields a data matrix $M$, shown in Figure 1 (right) for $\delta=0.3$ (the highest noise level)  [^2] Full version figure with an factorization result is reported in Appendix 6.7..
@@
Table 1 shows the mean error $\bar{\epsilon}$ and the standard error $s_{E}$ of the detected $k$ versus ground-truth $k^{*}=10$, across 30 generated datasets. We tested five popular detection algorithms: BCV [36], $BIC_{1}$ [53] [^3] Here $BIC_{1}(k):=\log\left(\|M-WQ^{T}\|_{F}^{2}\right)+k\frac{m+n}{mn}\log\left(\frac{mn}{m+n}\right)$, other versions yield similar results., CCC [22] and Dispersion [13]. For ICQF and $\ell_{1}$-NMF, BCV is the best detection scheme at all noise levels; $BIC_{2}$ performs well for low noise only. For the three common FA schemes, Horn’s PA [33] and MAP [57] are superior to $BIC_{2}$ [48], which aligns with empirical observations in [58, 59, 27]. ICQF with BCV outperforms $\ell_{1}$-NMF and FA at all noise levels.
@@ 4.2.1 Data
The 2001 Child Behavior Checklist (CBCL) is a general-purpose questionnaire covering different domains of psychopathology designed to screen and refer patients to pediatric psychiatry clinics, for a variety of diagnoses [31, 6, 7]. The referral is based either on raw answers on the questionnaire or syndrome-specific subscales derived from them. The checklist includes 113 questions, grouped into 8 syndrome subscales: Aggressive, Anxiety/Depressed, Attention, Rule Break, Social, Somatic, Thought, Withdrawn problems. Answers are scored on a three-point Likert scale (0=absent, 1=occurs sometimes, 2=occurs often) and the time frame for the responses is the past 6 months. We use the parent-reported CBCL responses.
@@ 4.2.2 Experimental setup
Baseline methods  Our first baseline method is $\ell_{1}$- regularized NMF ($\ell_{1}$-NMF) [15], as it also imposes non-negativity and sparsity constraints. As constructs (or questions) can be correlated, we rule out other NMF methods with orthogonality constraints. FA with promax rotation (FA-promax) [32] using minimum residual as estimation method is included because it is the most commonly used technique for analyzing questionnaires and extracting latent constructs. It is also a baseline familiar to the clinical community designing questionnaires. Finally, syndrome subscales are included since they are often used for diagnostic prediction in screening. To estimate the number of factors $k$, we use BCV for $\ell_{1}$-NMF and ICQF, and Horn’s parallel analysis for FA, the best approach for each method in the synthetic questionnaire experiments in Section 4.1.
Dataset splits  Within each dataset, we first split the participants into development and held-out sets with an $80/20$ ratio. The assignment is done using stratified sampling, to keep the distribution of confounds and diagnostic labels similar across both sets. Training and validation sets are derived from the development set, as explained in each experiment. All the quantitative results are obtained on the held-out set. To increase the robustness of our analysis, and obtain measures of uncertainty, we use different seeds to resample 30 dataset splits, and carry out experiments on each split. The reported results are obtained by averaging the results on the held-out set across all 30 splits.
Model training and inference  Let $W^{\text{set}}$ denote the participant factor matrix in ICQF or NMF, or the factor score in FA, with the superscript denoting the set. Similarly, let $Q$ denote the question loadings associated with a factor in each method. Model training will yield a ($W^{\text{train}}$, $Q$) for participants in the training set. Inference with the model will produce $W^{\text{validate}}$ and $W^{\text{held-out}}$ in validation and held-out sets, using the trained $Q$ and confounds $C^{\text{validate}},C^{\text{held-out}}$ (if applicable).
Figure 2: Heatmap of factor loadings $Q:=[{}^{R}\!Q,{}^{C}\!Q]$ from ICQF for factors proper, old/young and male/female confounds, and the implicit intercept (top) and loadings $Q$ from Factor Analysis with promax rotation (bottom). Abbreviated questions are listed at the bottom of each column. Questions are grouped by syndrome subscale; some factors are syndrome specific, while others bridge syndromes. For the enlarge version, see Appendix 6.8.
@@ 4.2.3 Experiment 1: qualitative comparison of ICQF with FA
Given the absence of ground-truth factorizations for this questionnaire, the qualitative assessment hinges on the relation of question loadings to the syndrome subscales used in clinical practice. While there were factors that loaded primarily in questions from one subscale, as expected, we were encouraged by finding others that grouped questions from multiple subscales, in ways that were deemed sensible co-occurrences by our clinical collaborators. As a further, sanity check, we inspected the loadings of confound Old (increasing age) and observe that they covered issues such as “Argues”, “Act Young”, “Swears” and “Alcohol”. The loadings of $Q$ also reveal the relative importance among questions in each estimated factor; subscales deem all questions equally important. Visualization of each latent factor and their corresponding themes are reported in Appendix 6.9. The themes are determined by our clinical collaborators, taking into consideration the weights assigned to each factor.
For comparison, Figure 2 (bottom) shows the loadings $Q$ from Factor Analysis with promax rotation. By means of parallel analysis, we have identified a value of $k=13$, which significantly exceeds the 8 syndrome subscales that were initially established during the development of the checklist. The absence of sparsity and non-negativity control also results in a matrix that is more densely populated with both positive and negative elements, in an arbitrary range. This can present challenges when attempting to interpret the loadings in conjunction with the factor matrix $W$, also without constraints. A detailed subjective evaluation showcasing the improvement of the propsoed ICQF against FA is reported in Appendix 6.9.
@@ 4.2.4 Experiment 2: preservation of diagnostic-related information
Our first quantitative metric to compare ICQF with baseline methods is the degree to which the low-dimensional factor representation of each participant (row of $W$) retains diagnostic information, across all 11 conditions we consider. Furthermore, this metric must be evaluated as a function of training sample size. As the sample size decreases, the regularization imposed by each method becomes more influential in determining the relationship between questions.
We evaluate this by creating training sets of different sizes from the development set (80, 40, 60, and 20 % of participants, with a fixed 20% as a validation set) and factorizing each of them with ICQF and the other methods. This yields a ${W}^{\text{train}}$, ${Q}^{\text{train}}$ for each combination of method and training set size, which is then used to infer factor scores $W^{\text{held-out}}_{\%}$ from the held-out set. The same held out-set is used for every method and dataset size being compared.
To estimate diagnostic prediction performance for each ${W}^{\text{train}}$, ${Q}^{\text{train}}$ factorization, we train a separate logistic regression model with $\ell_{2}$ regularization and balanced class weights from ${W}^{\text{train}}$ for each of the 11 diagnostic labels (i.e., 11 binary classification problems). The regularization strength is fine-tuned using ${W}^{\text{validate}}$, and prediction assessment is carried out on ${W}^{\text{held-out}}$ using the receiver operating characteristic (ROC) area under the curve (AUC) metric. The use of AUC is motivated from a clinical perspective, where clinicians often apply varying thresholds for detection depending on the aim of prediction, such as screening or intervention that incurs significant costs. We repeat this procedure in both CBCL-HBN and CBCL-ABCD data.
Figure 3 shows the trend and variability (95% confidence region) of the averaged AUCs of ICQF and the baseline methods using different dataset sizes (proportions of subjects), for HBN (left) and ABCD (right). In both HBN and ABCD, the ICQF outperforms other optimal baseline methods in maintaining high AUC scores across 11 conditions, and the difference in performance increases as the sample size decreases ($p\leq 0.01$, based on a one-side Wilcoxon signed rank test and adjusted using False Discovery Rate $\alpha=0.01$), except for $\ell_{1}$-NMF at $20\%$ in CBCL-HBN). Moreover, the factorization solutions obtained with ICQF are more stable in terms of the number of dimensions $k$ ($k=8$to 6 for ICQF, versus 8 $\to$ 3 for $\ell_{1}$-NMF and 13 $\to$ 18 for FA-promax in HBN; $k=7$to 7 for ICQF, versus 5 $\to$ 4 for $\ell_{1}$-NMF and 20 $\to$ 17 for FA-promax in ABCD). This is particularly noteworthy in comparison to $\ell_{1}$-NMF, as it indicates the extra bounded constraints on $W$ and the approximation matrix $M_{approx}$ makes BCV detect $k$ more consistently.
@@ 4.2.5 Experiment 3: quality of the factor loadings
Table 2: Top 2: Quality of $Q$ factor loadings at various training set sizes, within dataset. The values are the mean and standard deviation of Pearson correlation coefficients between best-matched $Q$ factors from the full dataset, and from decreasing size subsets of it. Bolded where ICQF is significantly better. Bottom: Agreement in $Q$ factor loadings between models estimated in CBCL in two independent datasets, measured in the same way.
@@
Our second quantitative metric to compare ICQF with baseline methods considers the change in quality of the factor loading matrix $Q$ as training sample size decreases, to examine the effect of regularization in constraining estimates. As before, we obtain a ${W}^{\text{train}}$, ${Q}^{\text{train}}$ for each combination of method and training set size. We then compare the loading matrix each size ($Q_{\%}$) with the one obtained on the full development dataset ($Q_{\text{full}}$). We do this by greedily matching each row from $Q_{\text{full}}$ with a row from $Q_{\%}$ by their Pearson correlation, and then computing the average correlation across pairs as the score. Given that a factorization learned on a smaller dataset may have fewer factors, we do this over the first $\min(k_{\text{full}},k_{\%}$) rows only. The first two rows of Table 2 reports this score for ICQF and the two baseline factorization methods, at each dataset size, on both CBCL-HBN and CBCL-ABCD datasets. ICQF outperforms the other methods at every dataset size ($p\leq 0.01$, based on a one-side Wilcoxon signed rank test and adjusted using False Discovery Rate $\alpha=0.01$), except for $\ell_{1}$-NMF at $20\%$ in CBCL-HBN.

### 17 Mar 2026 v2

@@ 3.2 Solving the optimization problem
The full proof of Proposition 3.2 is provided in Appendix D.
The two propositions, combined, show that our factorization can capture the true latent structure of the data, under the right conditions. The first is a linear combination of factors being a good approximation, which is the case for questionnaires. The second is having a robust estimator of the number of factors $k$, which we discuss next.
Choice of number of factors  For each $\beta$, we choose the number of factors $k$ using blockwise-cross-validation (BCV). Given a matrix $M$, for each possible $k$ considered, we permute the rows and columns of $M$. When confound variables $C$ are available, the row permutation can also be performed in a stratified manner, so as to preserve the distribution of $C$. The permuted matrix is then partitioned into $b_{r}\times b_{c}$ blocks, each of which is assigned into one of 10 folds. We then carry out a loop over folds where, for each fold, we 1) omit the blocks in that fold from the matrix, 2) factorize the remainder of the matrix, 3) impute the omitted blocks via matrix completion, and 4) compute the reconstruction error  [^1] A multiplicative correction factor is applied to the error if number of blocks in the last fold is less than in others. over omitted blocks. We choose $k$ with the lowest average error over folds. We compared this with other approaches for choosing $k$, for ICQF and other methods, over synthetic data, and report the results in Section 4.1.
@@ 4 Experiments and results
Across experiments, we compare ICQF with the methods most likely to be used in factorizing questionnaires. Our first baseline method is $\ell_{1}$- regularized NMF ($\ell_{1}$-NMF) Cichocki & Phan (2009), as it also imposes non-negativity and sparsity constraints. As constructs (or questions) can be correlated, we rule out other NMF methods with orthogonality constraints. FA with promax rotation (FA-promax) Hendrickson & White (1964) using minimum residual as estimation method is included because it is the most commonly used technique for analyzing questionnaires and extracting latent constructs. It is also a baseline familiar to the clinical community designing questionnaires. Both ICQF and $\ell_{1}$-NMF were initialized with NNDSVD Boutsidis & Gallopoulos (2008), and the same sparsity regularization strength ($\beta=1\mathrm{e}{-1}$) and stopping criterion (relative iteration convergence tolerance $\epsilon<1\mathrm{e}{-3})$ for fairness. The estimation method for FA was minimum residual.
@@ 4.1 Experiments on synthetic questionnaire data
Our first experiment was aimed at examining the effectiveness of BCV and other algorithms at estimating the number of latent factors when paired with the three factorization methods. To that effect we generated a realistic synthetic questionnaire with $k^{*}=10$ factors as ground truth. To generate a $200\times 10$ latent factor matrix $W$ (Figure 1 left), we first defined a matrix $D$, a binary matrix in which columns are correlated with a step-like pattern. Each column is present in isolation for 20 participants, and in tandem with another for 10 more, to synthesize correlation between factors. Then an entry of $W[i,j]$ is defined as
@@
Each factor had an associated loading vector – answer pattern – over 100 questions ($[0,100]$ range). The resulting $100\times 10$ loading matrix $Q$ , shown in Figure 1 (center), was defined to be
@@
We then created a noiseless data matrix $M_{clean}:=\min(0,\max(WQ^{T},100))$, and added noise by
@@
where $e$ follows a discrete probability distribution with $P(e=1)=\delta,P(e=0)=1-\delta$. This yielded a data matrix $M$, shown in Figure 1 (right) for $\delta=0.3$ (the highest noise level)  [^2] Full version figure with an factorization result is reported in Appendix G..
@@
Table 1 shows the mean error $\bar{\epsilon}$ and the standard error $s_{E}$ of the detected $k$ versus ground-truth $k^{*}=10$, across 30 generated datasets. We tested five popular detection algorithms: BCV Kanagal & Sindhwani (2010), $BIC_{1}$ Stoica & Selen (2004) [^3] Here $BIC_{1}(k):=\log\left(\|M-WQ^{T}\|_{F}^{2}\right)+k\frac{m+n}{mn}\log\left(\frac{mn}{m+n}\right)$, other versions yield similar results., CCC Fogel et al. (2007) and Dispersion Brunet et al. (2004). For ICQF and $\ell_{1}$-NMF, BCV was the best detection scheme at all noise levels; $BIC_{2}$ performed well for low noise only. For the three common FA schemes, Horn’s PA Horn (1965) and MAP Velicer (1976) were superior to $BIC_{2}$ Preacher et al. (2013), which aligns with empirical observations in Velicer et al. (2000); Watkins (2018); Goretzko et al. (2021). ICQF with BCV outperformed $\ell_{1}$-NMF and FA at all noise levels.
@@ 4.2.1 Data
The 2001 Child Behavior Checklist (CBCL) is a questionnaire covering different domains of psychopathology, designed to screen patients for referral to pediatric psychiatry clinics, for a variety of diagnoses Heflinger et al. (2000); Biederman et al. (2005; 2020).The checklist includes 113 questions, grouped into 8 syndrome subscales: Aggressive, Anxiety/Depressed, Attention, Rule Break, Social, Somatic, Thought, Withdrawn problems. The referral is typically based on scores obtained by adding answers on syndrome-specific subscales. Answers are scored on a three-point Likert scale (0=absent, 1=occurs sometimes, 2=occurs often) and the time frame for the responses is the past 6 months. We use the parent-reported CBCL responses.
@@ 4.2.2 Experimental setup
Baseline methods  As explained above, we use $\ell_{1}$-regularized NMF ($\ell_{1}$-NMF) Cichocki & Phan (2009) and factor analysis with promax rotation Hendrickson & White (1964) as baseline methods. We also include syndrome subscale scores, since they are often used for diagnostic prediction in screening. To estimate the number of factors $k$, we use BCV for $\ell_{1}$-NMF and ICQF, and Horn’s parallel analysis for FA, the best approach for each method in the synthetic questionnaire experiments in Section 4.1.
Dataset splits  Within each dataset, we first split the participants into development and held-out test sets with an $80/20$ ratio. The assignment is done using stratified sampling, to keep the distribution of confounds and diagnostic labels similar across both sets. Training and validation sets are derived from the development set, as explained in each experiment, and used to set all the parameters for testing. All the quantitative results are obtained on the held-out set. To increase the robustness of our analysis, and obtain measures of uncertainty, we use different seeds to resample 30 dataset splits, and carry out experiments on each split. The reported results are obtained by averaging the results on the held-out set across all 30 splits.
Model training and inference  Let $W^{\text{set}}$ denote the participant factor matrix in ICQF or NMF, or the analogous factor score matrix in FA, with the superscript denoting the dataset. Similarly, let $Q$ denote the question loadings associated with a factor in each method. Model training will yield a ($W^{\text{train}}$, $Q$) for participants in the training set. Inference with the model will produce $W^{\text{validate}}$ and $W^{\text{held-out}}$ in validation and held-out sets, using the trained $Q$ and confounds $C^{\text{validate}},C^{\text{held-out}}$ (if applicable).
Figure 2: Heatmap of factor loadings $Q:=[{}^{R}\!Q,{}^{C}\!Q]$ from ICQF for factors proper, old/young and male/female confounds, and the implicit intercept (top) and loadings $Q$ from Factor Analysis with promax rotation (bottom). Abbreviated questions are listed at the bottom of each column. Questions are grouped by syndrome subscale; some factors are syndrome specific, while others bridge syndromes. For the enlarge version, see Appendix H.
@@ 4.2.3 Experiment 1: qualitative comparison of ICQF with FA
Given the absence of ground-truth factorizations for this questionnaire, the qualitative assessment hinges on the relation of question loadings to the syndrome subscales used in clinical practice. While there were factors that loaded primarily in questions from one subscale, as expected, we were encouraged by finding others that grouped questions from multiple subscales, in ways that were deemed sensible by our clinical collaborators. As a further check, we inspected the loadings of confound Old (increasing age) and observe that they covered issues such as “Argues”, “Act Young”, “Swears” and “Alcohol”. The loadings of $Q$ also reveal the relative importance among questions in each estimated factor; for comparison, subscales would deem all questions equally important. Visualization of each latent factor and their corresponding themes and questions are reported in Appendix I. The themes were determined by our clinical collaborators, taking into consideration the weights each factor assigned across questions.
For comparison, Figure 2 (bottom) shows the loadings $Q$ from Factor Analysis with promax rotation. Using parallel analysis, we identified a value of $k=13$, which significantly exceeds the 8 syndrome subscales that were initially established during the development of the checklist. The absence of sparsity and non-negativity control also results in a matrix that is more densely populated with both positive and negative elements, in an arbitrary range. This can present challenges when attempting to interpret the loadings in conjunction with the factor matrix $W$, also without constraints. A detailed subjective evaluation by our clinical collaborators, showcasing the improvement of the proposed ICQF against FA, is given in Appendix I.
@@ 4.2.4 Experiment 2: preservation of diagnostic-related information
Our first quantitative metric to compare ICQF with baseline methods is the degree to which the low-dimensional factor representation of each participant (row of $W$) retains diagnostic information, across all 11 conditions we consider. Furthermore, this metric must be evaluated as a function of training sample size. As typical clinical studies have sample sizes an order of magnitude smaller than these, the regularization imposed by each method becomes more influential in determining the relationship between questions.
We evaluate this by creating training sets of different sizes from the development set (80, 40, 60, and 20 % of participants, with a fixed 20% as a validation set) and factorizing each of them with ICQF and the other methods. This yields a ${W}^{\text{train}}$, ${Q}^{\text{train}}$ for each combination of method and training set size, which is then used to infer factor scores $W^{\text{held-out}}_{\%}$ from the held-out set. The same held-out set is used for every method and dataset size being compared.
To estimate diagnostic prediction performance for each ${W}^{\text{train}}$, ${Q}^{\text{train}}$ factorization, we train a separate logistic regression model with $\ell_{2}$ regularization and balanced class weights from ${W}^{\text{train}}$ for each of the 11 diagnostic labels (i.e., 11 binary classification problems). The regularization strength is fine-tuned using ${W}^{\text{validate}}$, and prediction assessment is carried out on ${W}^{\text{held-out}}$ using the receiver operating characteristic (ROC) area under the curve (AUC) metric. The use of AUC is motivated from a clinical perspective, where clinicians often apply varying thresholds for detection depending on downstream use, such as screening versus intervention. We ran this procedure in both CBCL-HBN and CBCL-ABCD data.
Figure 3 shows the trend and variability (95% confidence region) of the average AUCs across 11 diagnostic problems, for ICQF and the baseline methods, using different dataset sizes (proportions of subjects), for HBN (left) and ABCD (right). In both HBN and ABCD, the ICQF outperforms other optimal baseline methods in maintaining high AUC scores across 11 conditions, and the difference in performance increases as the sample size decreases ($p\leq 0.01$, based on a one-side Wilcoxon signed rank test and adjusted using False Discovery Rate $\alpha=0.01$), except for $\ell_{1}$-NMF at $20\%$ in CBCL-HBN). Moreover, the factorization solutions obtained with ICQF are more stable in terms of the number of dimensions $k$ ($k=8$ $\to$ 6 for ICQF, versus 8 $\to$ 3 for $\ell_{1}$-NMF and 13 $\to$ 18 for FA-promax in HBN; $k=7$ $\to$ 7 for ICQF, versus 5 $\to$ 4 for $\ell_{1}$-NMF and 20 $\to$ 17 for FA-promax in ABCD). This is particularly noteworthy in comparison to $\ell_{1}$-NMF, as it indicates the extra bounded constraints on $W$ and the approximation matrix $M_{approx}$ makes BCV detect $k$ more consistently.
@@ 4.2.5 Experiment 3: quality of the factor loadings
Table 2: CBCL-HBN and CBCL-ABCD: Quality of $Q$ factor loadings at various training set sizes, within dataset. The values are the mean and standard deviation of Pearson correlation coefficients between best-matched $Q$ factors from the full dataset, and from decreasing size subsets of it. Bolded where ICQF is significantly better. CBCL-HBN $\leftrightarrow$ CBCL-ABCD: Agreement in $Q$ factor loadings between models estimated in CBCL in two independent datasets, measured in the same way.
@@
Our second quantitative metric to compare ICQF with baseline methods considers the change in quality of the factor loading matrix $Q$ as training sample size decreases, to examine the effect of regularization in constraining estimates. Ideally, a method should produce similar loadings across questions for each factor even as training datasets get smaller. As before, we obtain a ${W}^{\text{train}}$, ${Q}^{\text{train}}$ for each combination of method and training set size. We then compare the loading matrix at each size ($Q_{\%}$) with the one obtained on the full development dataset ($Q_{\text{full}}$). We do this by greedily matching each row from $Q_{\text{full}}$ with a row from $Q_{\%}$ by their Pearson correlation, and then computing the average correlation across pairs as the score. Given that a factorization learned on a smaller dataset may have fewer factors, we do this over the first $\min(k_{\text{full}},k_{\%}$) rows only. The first two rows of Table 2 reports this score for ICQF and the two baseline factorization methods, at each dataset size, on both CBCL-HBN and CBCL-ABCD datasets. ICQF outperforms the other methods at every dataset size ($p\leq 0.01$, based on a one-side Wilcoxon signed rank test and adjusted using False Discovery Rate $\alpha=0.01$), except for $\ell_{1}$-NMF at $20\%$ in CBCL-HBN.